Investment: $20,000 Annual Income: $1,280
x = T-Bills, 5% yeild y = T-Bonds, 7% yeild z = Corp Bonds, 9% yeild
amount invested in t-bills must be 2 times the amount invested in Corp Bills.
Find the amount in each investment.
I have two of the three equations:
x+y+z= 20,000
5x+7y+9z=128,000
What would the third equation be?

Answer by stanbon(57962)

Investment: $20,000 Annual Income: $1,280
x = T-Bills, 5% yeild y = T-Bonds, 7% yeild z = Corp Bonds, 9% yeild
amount invested in t-bills must be 2 times the amount invested in Corp Bills.
Find the amount in each investment.
I have two of the three equations:
x+y+z= 20,000
5x+7y+9z=128,000
What would the third equation be?
y = 2*z
================
Cheers,
Stan H.